A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week?
Answer by ankor@dixie-net.com(22740)

A girl scout troop sold cookies.
If the girls sold 5 more boxes the second week than they did the first,
and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies,
how many did they sell each week?
:
Let x = no. sold the 1st week
" the girls sold 5 more boxes the second week"
(x+5) = no. sold the 2nd week
" if they doubled the sales of the second week for the third week"
2(x+5) = no. sold the 3rd week
to sell a total of 431 boxes of cookies,?
x + (x+5) + 2(x+5) = 431
x + x + 5 + 2x + 10 = 431
4x + 15 = 431
4x = 431 - 15
4x = 416
x = 416/4
x = 104 sold the 1st week
then
104 + 5 = 109 sold the 2nd week
and
2(109) = 218 sold the 3rd week
:
:
Check: 104 + 109 + 218 = 431